OpenStudy (anonymous):

if someone can help me please respond here:

OpenStudy (amistre64):

whatcha need help with?

OpenStudy (anonymous):

how do i solve this system: 5x+2y =7 -x/2 + y/6 = 3/2 i have been trying to solve this but can't get it right, could you show me the steps!

OpenStudy (amistre64):

I did that one on your other post, do you see it?

OpenStudy (amistre64):

solve for y in the top: y = -5x/2 +7/2 plug it into the bottom equation: -x/2 + (-5x/2 +7/2)/6 = 3/2 Solve for x: -x/2 -5x/12 +7/12 = 3/2 -6x/12 -5x/12 +7/12 = 18/12 -6x -5x +7 =18 -11x = 11 x=-1 Now that we have a solid number for x; plug it into the equation we found for y: y = -5(-1)/2 +7/2 y= 5/2 + 7/2 y= 12/2 = 6 Double check for accuracy :)

OpenStudy (anonymous):

ok thanks i really appreciate it, i have one other... completing the square t^2 -9t_______ it says to first complete trinomial ____________________________ the then (_________________________)^2 bionomial squared...

OpenStudy (anonymous):

for the first question... i was taught to plug 0 in.... how did you get 7/2 ....

OpenStudy (amistre64):

Yes, to complete the square we need to find a term whose product adds up to (-9); which is just (-9/2)^2.

OpenStudy (amistre64):

To solve for y: subtract (5x) from both sides: 2y = -5x + 7 now divide everything by (2): (2/2)y = (-5/2)x + (7/2) y = (-5/2)x + (7/2)

OpenStudy (anonymous):

ok got that... and then this step confuses me Solve for x: -x/2 -5x/12 +7/12 = 3/2 -6x/12 -5x/12 +7/12 = 18/12

OpenStudy (amistre64):

(t-(9/2))^2 = (t-(9/2)) (t-(9/2)) = t^2 -9 + 81/4

OpenStudy (anonymous):

Solve for x: -x/2 -5x/12 +7/12 = 3/2 -6x/12 -5x/12 +7/12 = 18/12

OpenStudy (anonymous):

got me lost...

OpenStudy (anonymous):

it started out different...

OpenStudy (anonymous):

with the 6 on the bottoms and turned to tweleve why did you multiply

OpenStudy (amistre64):

To solve for x: we need to get all the fractions having the same denominators. The common denominator is 12: -x/2 needs to be multiplied by 6/6 to get a 12 on the bottom: -6x/12 -5x/12 already gots a 12 +7/12 already gots a 12 3/2 needs to be multiplied by 6/6 to get 12 on bottom: 18/12 Now because everything has a 12 on the bottom, we can toss the 12 out to clean up the equation: -6x -5x +7 = 18

OpenStudy (anonymous):

but what happens to the 6 on the bottom of the 2nd fraction...

OpenStudy (amistre64):

i see it: we solved for y to get: y = -5x/2 +7/2 The second equation has the term: y/6, so plug in what you found out about "y" to get: (-5x/2 +7/2)/6 when we divide the 6 thru we get: (-5x/2)/6 and (7/2)/6 the 6 actually multiplies to the bottom number, the 2. so that the denominators become 12. Does that make sense? I can try explaining it another way if you like

OpenStudy (anonymous):

i am writting it down sorry its taking me so long.. i really appreciate your help :)

OpenStudy (anonymous):

ok i understand that now, so then once i have those answers (-x/2) (-5x/2)/6 (7/2)/6 = 3/2 what am i doing to get twelve? becaues isn't (7/2)/6 technically .58

OpenStudy (amistre64):

let me try it this way, cause fractions are terrible in this format. top: (-5x)(1) +(7)(1) = -5x +7 bottom: (2) (6) (2)(6) = 12 12 How does that look? :)

OpenStudy (anonymous):

more confusing i have wrritting out fine... with the fractions over 6 but don't get what i am doing if i have 7/2 over 6 that equals .58

OpenStudy (amistre64):

To divide fractions; flip the bottom one over and multiply across. (7/2)/(6/1) is the same as (7/2)(1/6) = (7/12)

OpenStudy (amistre64):

gotta run; have fun :)

OpenStudy (anonymous):

thx